Smith-Kline Company maintains inventory records at selling prices as well as at cost. For 2021, the records indicate the following data:

($ in 000s)
Cost Retail
Beginning inventory $ 82 $ 127
Purchases 673 1,008
Freight-in on purchases 32
Purchase returns 1 2
Net markups 3
Net markdowns 8
Net sales 916

Use the retail method to approximate cost of ending inventory valued under Average cost method
Answer #1
Computation of the ending inventory under Average cost method
Cost ($ in 000S) Retail ($ in 000S)
Beginning inventory                   82.00                127.00
Purchases                 673.00              1,008.00
Freight- in on purchases                   32.00
Purchase returns                   (1.00)                  (2.00)
Net markups                       -                      3.00
Net markdowns                       -                    (8.00)
Goods available for sale                 786.00              1,128.00
Deduct: Net sales               (916.00)
Ending inventory, at retail                212.00
Average cost ratio = 786 / 1,128 = 69.68%
Ending inventory under Average cost method = 212 * 69.68% = $147.72